OpenStudy (jennychan12):

How to do Errors in Trapezoidal Rule and Simpson's Rule? See below for question.

OpenStudy (jennychan12):

\[\int\limits_{0}^{2} x^3dx\] when n = 4

OpenStudy (jennychan12):

@RadEn @Hero @agent0smith

OpenStudy (jennychan12):

anyone? @zepdrix

OpenStudy (jennychan12):

I know the formula is \[E \le \frac{ (b-a)^2 }{ 12n^2 } (\max \left| f"x \right|), a \le x \le b\] and I got 1/4 but the work here

OpenStudy (jennychan12):

anyone????

OpenStudy (agent0smith):

Are you sure you have the error formula correct? From looking it up, it looks like the (b-a) term should be to the power 3, so (b-a)^3 not (b-a)^2. I got 0.5 that way.

OpenStudy (jennychan12):

yeah i just got the error formula wrong...